# Effects of Near-Field Hydrodynamic Interactions on Bacterial Dynamics Near a Solid Surface

### v1

@@
Abstract
@@
Near-field hydrodynamic interactions between bacteria and no-slip solid surfaces are the main mechanism underlying surface entrapment of bacteria. In this study, we employ a chiral two-body model to simulate bacterial dynamics near the surface. The simulation results show that as bacteria approach the surface, their translational and rotational velocities, as well as their diffusion coefficients, decrease. Under the combination of near-field hydrodynamic interactions and DLVO forces, bacteria reach a stable fixed point in the phase plane and follow circular trajectories at this point. Notably, bacteria with left-handed helical flagella exhibit clockwise circular motion on the surface. During this process, as the stable height increases, the translational velocity parallel to the surface increases while the rotational velocity perpendicular to the surface decreases, collectively increasing the radius of curvature. Ultimately, our findings demonstrate that near-field hydrodynamic interactions significantly prolong the surface residence time of bacteria. Additionally, smaller stable heights further amplify this effect, resulting in longer residence times and enhanced surface entrapment.
preprint: AIP/123-QED
I Introduction
@@
The motility of bacteria near solid-liquid interfaces or on a solid surface plays a crucial role in fields such as biomedicine Tuson and Weibel (2013), and wastewater treatment Liu et al. (2013); Ghosh et al. (2019). Planktonic microorganisms, particularly flagellated bacteria, form biofilms through processes that include surface entrapment, migration, and colonization Monds and O’Toole (2009); Carniello et al. (2018); Krsmanovic et al. (2021). The resulting biofilms are closely associated with various infectious diseases Costerton, Stewart, and Greenberg (1999); Donlan (2001); Chang, Weeks, and Ducker (2018); Penesyan et al. (2021). Flagellated bacteria propel themselves to swim in fluid environments through the thrust generated by rotating their helical flagella Lauga (2016); Liu, Chen, and Xu (2025). Recent experiments using a combination of three-dimensional holographic microscopy and optical tweezers to track an individual Escherichia coli have revealed that surface entrapment of bacteria can be divided into three sequential stages: approach, reorientation, and surface swimming Bianchi, Saglimbeni, and Di Leonardo (2017); Bianchi et al. (2019). The physical properties of the solid surface significantly influence the dynamic behavior of bacteria Krsmanovic et al. (2021). In particular, on no-slip solid surfaces, near-field hydrodynamic interactions (HIs) between bacteria and the surface are the primary mechanism for surface entrapment Berke et al. (2008); Li et al. (2011); Wu, Hsiao, and Woon (2018); Zhang et al. (2021). However, HIs alone are insufficient to fully explain the circular motion of bacteria on surfaces. The Derjaguin–Landau–Verwey–Overbeek (DLVO) forces prevent bacteria from contacting the surface Sharma and Rao (2003); Hong et al. (2012); Liu, Jin, and An (2025). In the phase plane, the bacterial motion exhibits a stable fixed point under the combination of near-field HIs and DLVO forces Frymier et al. (1995); Li, Tam, and Tang (2008); Liu, Jin, and An (2025).
@@
Bacteria undergoing Brownian motion near a no-slip solid surface exhibit various patterns of motility, including surface entrapment Vissers et al. (2018); Berne et al. (2018), circular motion Li, Tam, and Tang (2008); Utada et al. (2014), and escape from surfaces Junot et al. (2022). Both experiments and numerical simulations have demonstrated that the translational and rotational velocities, the radius of curvature of the circular motion, and the surface residence time of bacteria are mainly influenced by near-field HIs between bacteria and the surface Lauga et al. (2006); Junot et al. (2022), Brownian motion Li, Tam, and Tang (2008), surface properties Khalid et al. (2020); Zheng et al. (2021); Krsmanovic et al. (2021), and bacterial morphology Di Leonardo et al. (2011); Tokárová et al. (2021). The vertical translational velocity of bacteria is reduced as their distance to the surface decreases due to the near-field HIs Lauga et al. (2006); Liu, Chen, and Zhang (2025). The shear flow induced by a no-slip solid surface exerts a torque on the cell body that tends to orient the bacteria toward the surface Petroff, Wu, and Libchaber (2015); Bianchi, Saglimbeni, and Di Leonardo (2017). In contrast, the drag force acting on the flagella drives the bacteria to orient parallel to the surface Sipos et al. (2015). Consequently, bacteria with shorter flagella often orient perpendicularly to the surface ("nose down") Petroff, Wu, and Libchaber (2015), while those with longer flagella swim parallel to the surface Petroff and Libchaber (2018); Das and Lauga (2019); Liu, Jin, and An (2025). Moreover, experiments indicate that the surface suppresses bacterial tumbling, preventing the escape of bacteria from the solid surface Molaei et al. (2014). The surface residence time of bacteria swimming in the surface region can exceed one hundred seconds Junot et al. (2022). Random forces and torques are generally considered the primary mechanism that allows bacteria to escape surface entrapment Schaar, Zöttl, and Stark (2015); Militaru et al. (2021), and longer flagella have been reported to facilitate such escape Sipos et al. (2015). These observations highlight the important role of near-field HIs of bacteria in bacterial dynamics near solid surfaces. However, the detailed role of near-field HIs in these orientation and escape processes remains poorly understood, motivating the present study.
@@
The goal of this article is to investigate how near-field HIs between bacteria and a no-slip solid surface influence bacterial velocities, the radius of curvature of circular motion, and their surface residence time. In our numerical simulations, bacteria are represented by a chiral two-body model Di Leonardo et al. (2011); Dvoriashyna and Lauga (2021); Liu et al. (2025). The $6\times 6$ resistance matrix of flagella is derived using resistive force theory (RFT) Gray and Hancock (1955); Chwang and Wu (1975); Johnson and Brokaw (1979), which incorporates flagellar morphology and chirality. The resistance matrix between the bacterial cell body and the surface is calculated using the method proposed by Dunstan et al., which accounts for near- and far-field HIs Dunstan et al. (2012).
@@
This article is organized as follows. The numerical simulation model and methods are presented in Sec. II, where we employ a chiral two-body model to represent the bacterium and to account for near-field HIs. Numerical simulations without Brownian motion and with Brownian motion are performed in Secs. III. A and III. B, respectively. The main conclusions are summarized in Sec. IV.
II Simulation Model
@@
Figure 1: Schematic diagram of a bacterium model with a rigid helical flagellum and a spherical cell body swimming near a no-slip solid surface. $\mathbf{r}_{b}$ and $\mathbf{r}_{t}$ are the center positions of the cell body and flagellum, respectively. The closest distance between the cell body and solid surface is $h$, and the inclination angle is $\Psi$. The surface is located on the xyxy-plane.
@@
A singly flagellated bacterium is modeled as a single left-handed helical flagellum attached to a spherical cell body. The closest distance between the cell body and the no-slip solid surface is $h$, as shown in Fig. 1. The surface is located on the $xy$-plane, and the connection point between the cell body and the flagellar axis is denoted as $\mathbf{r}_{0}$. The helical flagellum is modeled as a rigid helix with filament radius $a$, helix radius $R$, pitch $\lambda$, pitch angle $\theta$, axial length $L$, and contour length $\Lambda=L/\cos\theta$, where $\tan\theta=2\pi R/\lambda$. The centerline positions of the left-handed helical flagellum are given by:
@@
where $l\in[0,\Lambda]$ is the arc length along the centerline, $k=2\pi/\lambda$ is the wave number, and $\phi=kl\cos\theta$ is the phase. The time-dependent material frame at the connection point $\mathbf{r}_{0}$ is denoted as $\{\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{D}_{3}\}$. This is a set of orthonormal vectors, with $\mathbf{D}_{1}$ being the unit vector along the flagellar axis. The radius of the spherical cell body is $R_{b}$. The bacterium can be represented by a chiral two-body model, with the centers of the cell body and the flagellum denoted as $\mathbf{r}_{b}$ and $\mathbf{r}_{t}$, respectively. The distance between the centers of the cell body and the flagellum is $d=L/2+R_{b}$. The points $\mathbf{r}_{0}$ and $\mathbf{r}_{b}$ satisfy the relationship: $\mathbf{r}_{0}=\mathbf{r}_{b}-R_{b}\cdot\mathbf{D}_{1}$.
Table 1: Geometric parameters and corresponding values of a bacterium system.
Symbol Parameters Values
$\mu$ Dynamic viscosity $1.0$ $\mu\mathrm{g}\mathrm{/}\mathrm{(}\mu\mathrm{m}\cdot\mathrm{s}\mathrm{)}$
$R_{b}$ Radius of cell body $1.0$ $\mu\mathrm{m}$
$a$ Filament radius $0.01$ $\mu\mathrm{m}$
$R$ Helix radius $0.25$ $\mu\mathrm{m}$
$\theta$ Pitch angle $\pi/5$
$\Lambda$ Contour length $7.5$ $\mu\mathrm{m}$
$f$ Rotation frequency of motor $100$ $\mathrm{H}\mathrm{z}$
$H$ Hamaker constant $10^{-21}$ $\mathrm{J}$
$\zeta_{1}$ Zeta potential of cell body 20.0-20.0 $\mathrm{m}\mathrm{V}$
$\zeta_{2}$ Zeta potential of surface 20.0-20.0 $\mathrm{m}\mathrm{V}$
HIs between the cell body and the flagellum Lauga et al. (2006), and those between the flagellum and the surface, are neglected. The rigid helical flagellum is represented by a chiral body model, with its $6\times 6$ resistance matrix obtained by integrating the flagellar centerline using RFT and expressed as Liu et al. (2025)
@@
where the three submatrices are given by Di Leonardo et al. (2011); Dvoriashyna and Lauga (2021); Liu et al. (2025):
@@
where $\mathbf{e}_{a}=\mathbf{D}_{1}$ is the unit vector that indicates the direction of the flagellar axis. The specific expressions of these matrix elements are detailed in Appendix A. The resistance matrix for a spherical cell body near a solid surface can be written as Dunstan et al. (2012); Liu, Jin, and An (2025)
@@
The scalar functions $\bar{Y}_{\parallel}^{A}$, $\bar{Y}_{\perp}^{A}$, $\bar{Y}^{B}$, $\bar{Y}_{\parallel}^{C}$, and $\bar{Y}_{\perp}^{C}$ of the resistance matrix of the cell body are plotted in Fig. 2. It is evident that these functions are strongly dependent on the reduced height $h/R_{b}$. In particular, the function $Y_{\perp}^{A}$ is approximately $10$ times larger than its value in the bulk fluid at $h/R_{b}=0.1$, and diverges as $h/R_{b}$ approaches 0. The bacterial diffusion coefficient, $D=k_{B}T/\mathcal{R}$, where $k_{B}$ is the Boltzmann constant, $T$ is the absolute temperature and $\mathcal{R}$ is the resistance matrix. As bacteria approach the surface, their resistance increases, resulting in a decrease in their diffusion coefficient and effectively suppressing their mobility. The functions $Y_{\parallel}^{A}$, $Y_{\parallel}^{C}$, and $Y_{\perp}^{C}$ also increase as $h$ decreases, indicating that the bacterial velocities decrease closer to the surface. The function $Y^{B}$ characterizes the hydrodynamic coupling between rotational velocity and force, which can result in "nose down" configurations Das et al. (2015); Sipos et al. (2015); Petroff, Wu, and Libchaber (2015); Bianchi et al. (2019); Liu, Chen, and Zhang (2025).
Figure 2: The scalar functions (a) $\bar{Y}_{\parallel}^{A}$, $\bar{Y}_{\perp}^{A}$, (b) $\bar{Y}^{B}$, (c) $\bar{Y}_{\parallel}^{C}$ and $\bar{Y}_{\perp}^{C}$ of the resistance matrix as a function of reduced height $h/R_{b}$. (d) DLVO forces of a sphere interacting with a solid surface as a function of the closest distance $h$ for different Debye lengths.
In addition to the HIs, the DLVO force is also present between the cell body and the surface. The DLVO force comprises the sum of the van der Waals (vdW) force and the electrostatic (ele) force Sharma and Rao (2003); Hong et al. (2012); Liu, Jin, and An (2025):
@@
where $h$ is the closest distance between the cell body and the surface. The expressions for the van der Waals (vdW) force and the electrostatic (ele) force are given by Sharma and Rao (2003); Hong et al. (2012); Liu, Jin, and An (2025):
@@
where $\varepsilon=6.933\times 10^{-10}$ $\mathrm{C}^{2}\mathrm{/}\mathrm{(}\mathrm{N}\cdot\mathrm{m}^{2}\mathrm{)}$ is the permittivity of water at room temperature $T=298$ $\mathrm{K}$, $H$ is the Hamaker constant, $\kappa$ is the inverse Debye length, and $\zeta_{1}$ and $\zeta_{2}$ are the zeta potentials of the cell body and the solid surface, respectively.
[…]
The instantaneous translational and rotational velocities of the center of the cell body are $\mathbf{U}_{b}$ and $\mathbf{W}_{b}$, respectively, while those of the flagellar center are
@@
The force and torque balance equations of a bacterium that swims near a solid surface are
@@
where $\mathbf{F}_{b}^{h}$, $\mathbf{T}_{b}^{h}$, $\mathbf{F}_{t}^{h}$, and $\mathbf{T}_{t}^{h}$ are the forces and torques exerted on the fluid by the cell body and the flagellum, respectively. It should be noted that a negative sign is added to represent the forces and torques exerted by the fluid on the cell body and flagellum. $\mathbf{F}_{D}$ is the DLVO force between the cell body and the surface. $\mathbf{F}_{b}^{B}$, $\mathbf{T}_{b}^{B}$, $\mathbf{F}_{t}^{B}$, and $\mathbf{T}_{t}^{B}$ are the random forces and torques on the cell body and flagellum, respectively. The expressions for these random forces are as follows:
@@
When a bacterium is submerged in a viscous fluid, its motion is governed by the following equation Lauga et al. (2006); Lauga and Powers (2009); Liu, Jin, and An (2025):
@@
where $\mathbf{0}$ is a $6\times 6$ zero matrix. The center of the cell body is updated by Liu, Chen, and Xu (2025):
@@
and the material frame $\{\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{D}_{3}\}$ is updated as follows Liu, Chen, and Xu (2025):
@@
III Results
III.1 Simulation Without Brownian Motion

### v2

# Baopi Liu,a,b Lu Chen,c and Haiqin Wang d

@@
[Uncaptioned image] Effects of Near-Field Hydrodynamic Interactions on Bacterial Dynamics Near a Solid Surface
Baopi Liu,a,b Lu Chen,c and Haiqin Wang d
@@
[Uncaptioned image] Near-field hydrodynamic interactions between bacteria and no-slip solid surfaces are the main mechanism underlying surface entrapment of bacteria. In this study, we employ a chiral two-body model to simulate bacterial dynamics near the surface. The simulation results show that as bacteria approach the surface, their translational velocities and diffusion coefficients decrease. Under the combination of near-field hydrodynamic interactions and DLVO forces, bacteria reach a stable fixed point in the phase plane and follow circular trajectories at this point. In particular, bacteria with left-handed helical flagella exhibit clockwise circular motion on the surface. During this process, as the stable height increases, the near-field hydrodynamic interactions weaken. Consequently, the translational velocity of the bacteria parallel to the surface increases while the rotational velocity perpendicular to the surface decreases, collectively increasing the radius of curvature. Ultimately, our findings demonstrate that near-field hydrodynamic interactions significantly prolong the surface residence time of bacteria. Additionally, smaller stable heights further amplify this effect, resulting in longer residence times and enhanced surface entrapment.
footnotetext: E-mail: bpliu@mail.bnu.edu.cnfootnotetext: a School of Advanced Interdisciplinary Studies, Ningxia University, Zhongwei, 755000, Chinafootnotetext: b Complex Systems Division, Beijing Computational Science Research Center, Beijing 100193, Chinafootnotetext: c College of Physics, Changchun Normal University, Changchun, Jilin 130032, Chinafootnotetext: d Department of Chemistry, Harvey Mudd College, Claremont, California 91711, USA
1 Introduction
@@
The motility of bacteria near solid-liquid interfaces or on a solid surface plays a crucial role in fields such as biomedicine 49, and wastewater treatment 34, 18. Planktonic microorganisms, particularly flagellated bacteria, form biofilms through processes that include surface entrapment, migration, and colonization 38, 6, 24. The resulting biofilms are closely associated with various infectious diseases 9, 13, 7, 39. Flagellated bacteria propel themselves to swim in fluid environments through the thrust generated by rotating their helical flagella 27, 30. Recent experiments using a combination of three-dimensional holographic microscopy and optical tweezers to track an individual E. coli have revealed that surface entrapment of bacteria can be divided into three sequential stages: approach, reorientation, and surface swimming 4, 5. The physical properties of the solid surface significantly influence the dynamic behavior of bacteria 24. In particular, on no-slip solid surfaces, near-field hydrodynamic interactions (HIs) between bacteria and the surface are the primary mechanism for surface entrapment 2, 28, 53, 54. However, HIs alone are insufficient to fully explain the circular motion of bacteria on surfaces. The Derjaguin–Landau–Verwey–Overbeek (DLVO) forces prevent bacteria from contacting the surface 44, 20, 32. In the phase plane, the bacterial motion exhibits a stable fixed point under the combination of near-field HIs and DLVO forces 17, 29, 32.
@@
Bacteria undergoing Brownian motion near a no-slip solid surface exhibit various patterns of motility, including surface entrapment 52, 3, circular motion 29, 50, and escape from surfaces 22. Both experiments and numerical simulations have demonstrated that the translational and rotational velocities, the radius of curvature of the circular motion, and the surface residence time of bacteria are mainly influenced by near-field HIs between bacteria and surfaces 25, 22, Brownian motion 29, surface properties 23, 55, 24, and bacterial morphology 12, 48. The vertical translational velocity of bacteria is reduced as their distance to the surface decreases due to the near-field HIs 25, 31. The shear flow induced by a no-slip solid surface exerts a torque on the cell body that tends to orient the bacteria toward the surface 40, 4. In contrast, the drag force acting on the flagella drives the bacteria to orient parallel to the surface 47. Consequently, bacteria with shorter flagella often orient perpendicularly to the surface ("nose down") 40, representing the surface-bound state where the flagellum points into the surface10. However, those with longer flagella tend to swim parallel to the surface, corresponding to the circular-swimming state 41, 10, 32. Moreover, experiments indicate that the surface suppresses bacterial tumbling, preventing the escape of bacteria from the solid surface 37. Random forces and torques are generally considered the primary mechanism that allows bacteria to escape surface entrapment 43, 36, and longer flagella have been reported to facilitate such escape 47. Experiments show that the surface residence time of bacteria that swim in the surface region can exceed one hundred seconds 22. These observations highlight the important role of near-field HIs of bacteria in bacterial dynamics near solid surfaces. However, the detailed role of near-field HIs in bacterial entrapment and escape processes remains poorly understood, motivating the present study.
@@
The goal of this article is to investigate how the near-field HIs between bacteria and a no-slip solid surface influence bacterial velocities, the radius of curvature of circular motion, and their surface residence time. In our numerical simulations, bacteria are represented by a chiral two-body model 12, 15, 33. The $6\times 6$ resistance matrix of flagella is derived using resistive force theory (RFT) 19, 8, 21, which includes flagellar morphology and chirality. The resistance matrix between the bacterial cell body and the surface is calculated using the method proposed by Dunstan et al., which accounts for near- and far-field HIs 14.
@@
This article is organized as follows. The numerical simulation model and methods are presented in Sec. 2, where we employ a chiral two-body model to represent the bacterium and to account for near-field HIs. Numerical simulations without Brownian motion and with Brownian motion are performed in Secs. 3.1 and 3.2, respectively.
2 Simulation Model
@@
Fig. 1: Schematic diagram of a bacterium model with a rigid helical flagellum and a spherical cell body swimming near a no-slip solid surface. $\mathbf{r}_{b}$ and $\mathbf{r}_{t}$ are the center positions of the cell body and flagellum, respectively. The closest distance between the cell body and solid surface is $h$, and the inclination angle is $\Psi$. The surface is located on the xyxy-plane.
@@
A single flagellated bacterium is modeled as a single left-handed helical flagellum attached to a spherical cell body, which is used as a simplified model for E. coli 25, 29, 47, 40. The closest distance between the cell body and the no-slip solid surface is $h$, as shown in Fig. 1. The surface is located on the $xy$-plane, and the connection point between the cell body and the flagellar axis is denoted as $\mathbf{r}_{0}$. The helical flagellum is modeled as a rigid helix with filament radius $a$, helix radius $R$, pitch $\lambda$, pitch angle $\theta$, axial length $L$, and contour length $\Lambda=L/\cos\theta$, where $\tan\theta=2\pi R/\lambda$. The centerline positions of the left-handed helical flagellum are given by 42, 16, 30:
@@
where $l\in[0,\Lambda]$ is the arc length along the centerline, $k=2\pi/\lambda$ is the wave number, and $\phi=kl\cos\theta$ is the phase. The time-dependent material frame at the connection point $\mathbf{r}_{0}$ is denoted as $\{\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{D}_{3}\}$. This is a set of orthonormal vectors, with $\mathbf{D}_{1}$ being the unit vector along the flagellar axis. The radius of the spherical cell body is $R_{b}$. The bacterium can be represented by a chiral two-body model, with the centers of the cell body and the flagellum denoted as $\mathbf{r}_{b}$ and $\mathbf{r}_{t}$, respectively. This model can be effectively used to simulate the Brownian motion of E. coli 33. The distance between the centers of the cell body and the flagellum is $d=L/2+R_{b}$. The points $\mathbf{r}_{0}$ and $\mathbf{r}_{b}$ satisfy the relationship: $\mathbf{r}_{0}=\mathbf{r}_{b}-R_{b}\cdot\mathbf{D}_{1}$.
Table 1: Geometric parameters and corresponding values of a bacterium system 25, 29, 32
. Symbol Parameters Values $\mu$ Dynamic viscosity $1.0$ $\mu\mathrm{g}\mathrm{/}\mathrm{(}\mu\mathrm{m}\cdot\mathrm{s}\mathrm{)}$ $R_{b}$ Radius of cell body $1.0$ $\mu\mathrm{m}$ $a$ Filament radius $0.01$ $\mu\mathrm{m}$ $R$ Helix radius $0.25$ $\mu\mathrm{m}$ $\theta$ Pitch angle $\pi/5$ $\Lambda$ Contour length $7.5$ $\mu\mathrm{m}$ $f$ Rotation frequency of motor $100$ $\mathrm{H}\mathrm{z}$ $H$ Hamaker constant $10^{-21}$ $\mathrm{J}$ $\zeta_{1}$ Zeta potential of cell body 20.0-20.0 $\mathrm{m}\mathrm{V}$ $\zeta_{2}$ Zeta potential of surface 20.0-20.0 $\mathrm{m}\mathrm{V}$
HIs between the cell body and flagellum, and between the flagellum and the surface, are neglected. HIs between the cell body and the flagellum contribute only a small error to the final results, because the presence of a nearby surface leads to spatially localized flow fields that decay as rapidly as a Stokeslet-dipole 25. HIs between the flagellum and surface become significant only at separations less than the filament radius ($a=10$ $\mathrm{n}\mathrm{m}$), a separation that only a small part of flagella can achieve. 29. The rigid helical flagellum is represented by a chiral body model, with its $6\times 6$ resistance matrix obtained by integrating the flagellar centerline using RFT and expressed as 33
@@
where the three submatrices are given by 12, 15, 33:
@@
where $\mathbf{e}_{a}=\mathbf{D}_{1}$ is the unit vector that indicates the direction of the flagellar axis. The specific expressions of these matrix elements are detailed in Appendix A. The resistance matrix for a spherical cell body near a solid surface can be written as 14, 32
@@
Fig. 2: The scalar functions (a) $\bar{Y}_{\parallel}^{A}$, $\bar{Y}_{\perp}^{A}$, (b) $\bar{Y}^{B}$, (c) $\bar{Y}_{\parallel}^{C}$ and $\bar{Y}_{\perp}^{C}$ of the resistance matrix as a function of reduced height $h/R_{b}$. (d) DLVO forces of a sphere interacting with a solid surface as a function of the closest distance $h$ for different Debye lengths.
The scalar functions $\bar{Y}_{\parallel}^{A}$, $\bar{Y}_{\perp}^{A}$, $\bar{Y}^{B}$, $\bar{Y}_{\parallel}^{C}$, and $\bar{Y}_{\perp}^{C}$ of the resistance matrix of the cell body are plotted in Fig. 2. It is evident that these functions are strongly dependent on the reduced height $h/R_{b}$. In particular, the function $Y_{\perp}^{A}$ at $h/R_{b}=0.1$ is approximately $10$ times larger than its value in the bulk fluid ($h/R_{b}\to\infty$), and diverges as $h/R_{b}$ approaches 0. The bacterial diffusion coefficient, $D=k_{B}T/\mathcal{R}$, where $k_{B}$ is the Boltzmann constant, $T$ is the absolute temperature and $\mathcal{R}$ is the resistance matrix. As bacteria approach the surface, their resistance increases, resulting in a decrease in their diffusion coefficient and effectively suppressing their motility, as shown in Appendix C. The functions $Y_{\parallel}^{A}$, $Y_{\parallel}^{C}$, and $Y_{\perp}^{C}$ also increase as $h$ decreases, indicating that the bacterial velocities decrease closer to the surface. The function $Y^{B}$ characterizes the hydrodynamic coupling between rotational velocity and force, which can result in "nose down" configurations 11, 47, 40, 5, 31.
In addition to the HIs, the DLVO force is also present between the cell body and the surface. The DLVO force comprises the sum of van der Waals (vdW) force and electrostatic (ele) force 44, 20, 32:
@@
where $h$ is the closest distance between the cell body and the surface. The expressions for the van der Waals (vdW) force and the electrostatic (ele) force are given by 44, 20, 32:
@@
where $\varepsilon=6.933\times 10^{-10}$ $\mathrm{C}^{2}\mathrm{/}\mathrm{(}\mathrm{N}\cdot\mathrm{m}^{2}\mathrm{)}$ is the permittivity of water at room temperature $T=298$ $\mathrm{K}$, $H$ is the Hamaker constant, and $\zeta_{1}$ and $\zeta_{2}$ are the zeta potentials of the cell body and the solid surface, respectively. The Debye length, $\kappa^{-1}$, is given by $\kappa^{-1}=0.304/\sqrt{I}$$\mathrm{n}\mathrm{m}$, where $I$ is the ionic strength (in $\mathrm{m}\mathrm{o}\mathrm{l}\mathrm{/}\mathrm{L}$) 32.
In the present paper, we set $H=10^{-21}$ $\mathrm{J}$ and the zeta potentials as $\zeta_{1}=\zeta_{2}=-20$ $\mathrm{m}\mathrm{V}$ 51, 29. Fig. 2(d) illustrates the DLVO forces between a sphere of radius $R_{b}=1$ $\mu\mathrm{m}$ and a solid surface for different Debye lengths. As the sphere approaches the surface, the DLVO forces increase rapidly with decreasing separation distance $h$, which can effectively prevent bacteria from contacting the surface during the Brownian motion. Bacteria reach a stable fixed point $\{h^{*},\Psi^{*}\}$ in the phase plane $\{h,\Psi\}$ under the combination of near-field HIs and DLVO forces 32. HIs are long-range forces, yet their influence becomes particularly significant at separations less than the particle radius, while DLVO forces are short-range forces, with an effective distance set at $100$ $\mathrm{n}\mathrm{m}$. The angle of inclination $\Psi$, defined as the angle between the flagellar axis and the $z$-axis, characterizes the orientation of the bacteria. Specifically, bacteria tend to exhibit circular motion on the surface at this stable fixed point 17.
The instantaneous translational and rotational velocities of the center of the cell body are $\mathbf{U}_{b}=\{U_{x},U_{y},U_{z}\}$ and $\mathbf{W}_{b}=\{W_{x},W_{y},W_{z}\}$, respectively, while those of the flagellar center are
@@
The force and torque balance equations governing the motion of a bacterium swimming near a solid surface are given by:
@@
where $\mathbf{F}_{b}^{h}$, $\mathbf{T}_{b}^{h}$, $\mathbf{F}_{t}^{h}$, and $\mathbf{T}_{t}^{h}$ are the forces and torques exerted on the fluid by the cell body and the flagellum, respectively. $\mathbf{F}_{D}$ is the DLVO force between the cell body and the solid surface. $\mathbf{F}_{b}^{B}$, $\mathbf{T}_{b}^{B}$, $\mathbf{F}_{t}^{B}$, and $\mathbf{T}_{t}^{B}$ represent the random forces and torques on the cell body and flagellum, respectively. The expressions for these random forces are as follows 29, 33:
@@
When a bacterium is submerged in a viscous fluid, its motion is governed by the following equation 25, 26, 32:
@@
where $\mathbf{0}$ is a $6\times 6$ zero matrix. The center of the cell body is updated by 30:
@@
and the material frame $\{\mathbf{D}_{1},\mathbf{D}_{2},\mathbf{D}_{3}\}$ is updated as follows 30:
@@
3 Results
3.1 Surface Swimming Dependent on Near-Field HIs
We first qualitatively and mechanistically examine the effect of near-field HIs on bacterial motility near a no-slip solid surface, excluding Brownian motion. When considering only near-field HIs, the translational and rotational velocities of the cell body as it gradually approaches the surface perpendicularly are:
$\begin{split}&U_{z}=\frac{Y_{\perp}^{C}X_{\parallel}^{B}}{(X_{\parallel}^{B})^{2}-(X_{\parallel}^{A}+Y_{\perp}^{A})(X_{\parallel}^{C}+Y_{\perp}^{C})}W_{0},\\ &W_{z}=\frac{X_{\parallel}^{A}X_{\parallel}^{C}-(X_{\parallel}^{B})^{2}+Y_{\perp}^{A}X_{\parallel}^{C}}{(X_{\parallel}^{B})^{2}-(X_{\parallel}^{A}+Y_{\perp}^{A})(X_{\parallel}^{C}+Y_{\perp}^{C})}W_{0}.\end{split}$ (14)